# BayesianVLA: Bayesian Decomposition of Vision Language Action Models via Latent Action Queries

### v3

@@ Abstract
Vision-Language-Action (VLA) models have shown promise in robot manipulation but often struggle to generalize to new instructions or complex multi-task scenarios. We identify a critical pathology in current training paradigms where goal-driven data collection creates a dataset bias. In such datasets, language instructions are highly predictable from visual observations alone, causing the conditional mutual information between instructions and actions to vanish, a phenomenon we term Information Collapse. Consequently, models degenerate into vision-only policies that ignore language constraints and fail in out-of-distribution (OOD) settings. To address this, we propose BayesianVLA, a novel framework that enforces instruction following via Bayesian decomposition. By introducing learnable Latent Action Queries, we construct a dual-branch architecture to estimate both a vision-only prior $p(a\mid v)$ and a language-conditioned posterior $\pi(a\mid v,\ell)$. We then optimize the policy to maximize the conditional Pointwise Mutual Information (PMI) between actions and instructions. This objective effectively penalizes the vision shortcut and rewards actions that explicitly explain the language command. Without requiring new data, BayesianVLA significantly improves generalization. Extensive experiments across on SimplerEnv and RoboCasa demonstrate substantial gains, including an 11.3% improvement on the challenging OOD SimplerEnv benchmark, validating the ability of our approach to robustly ground language in action.
BayesianVLA: Bayesian Decomposition of Vision Language Action Models via Latent Action Queries
Shijie Lian1,2,thanks: Equal contribution Bin Yu2,4,11footnotemark: 1 Xiaopeng Lin2,5,11footnotemark: 1 Laurence T. Yang6,1,thanks: Corresponding author Zhaolong Shen2,7 Changti Wu2,8 Yuzhuo Miao1,2 Cong Huang2,3 Kai Chen2,3,9,22footnotemark: 2 1HUST 2ZGCA 3ZGCI 4HIT 5HKUST(GZ) 6ZZU 7BUAA 8ECNU 9DeepCybo
@@ 1 Introduction
While demonstrating strong performance in in-distribution settings, current VLA models still face challenges in generalizing to novel instructions or complex multi-task scenarios, particularly in out-of-distribution (OOD) environments (Xu et al., 2025). This limitation is especially pronounced during post-training, where fine-tuning on narrow, task-specific datasets can lead to catastrophic forgetting of the VLM’s general capabilities and impair its ability to generalize to new tasks. We hypothesize that this fragility is exacerbated by a prevalent bias in current robotic datasets. Most robotic datasets are collected in a goal-driven manner, where a human operator performs a specific task repeatedly in a fixed scene. In such datasets, the mapping from visual scene $v$ to language instruction $\ell$ is nearly injective: seeing a cabinet in the scene almost invariably implies the task “open the cabinet,” while seeing a bottle implies “pick up the bottle.” This deterministic coupling results in a sharp conditional distribution $p(\ell\mid v)$.
@@
In other words, the model effectively ignores the language instruction, learning a “vision shortcut” that fails whenever the task is ambiguous or the environment changes.
To address this, we propose BayesianVLA, a novel framework that explicitly enforces instruction following via Bayesian decomposition. Our key insight is to maximize the conditional Pointwise Mutual Information (PMI) between actions and instructions, which is equivalent to maximizing the log-likelihood ratio (LLR): $\log p(\ell\mid a,v)-\log p(\ell\mid v)$. This objective penalizes the vision shortcut by requiring the action $a$ to provide additional information about $\ell$ that cannot be inferred from $v$ alone.
We instantiate this framework by introducing Latent Action Queries—a set of learnable tokens injected into the VLM. These queries serve a dual purpose: they act as a bottleneck to extract action-relevant features for a downstream Diffusion Transformer (DiT) policy, and they enable a dual-branch training strategy. In the Priori Branch, queries attend only to vision to learn $p(a\mid v)$; in the Posteriori Branch, they attend to both vision and language to learn $\pi(a\mid v,\ell)$. By optimizing the LLR between these branches, BayesianVLA learns to ground language robustly without requiring new data.
@@
    We propose BayesianVLA, a method that leverages Latent Action Queries and a dual-branch Bayesian objective to recover the true language-conditioned policy from biased data.
@@
    We demonstrate that BayesianVLA achieves state-of-the-art performance on SimplerEnv and RoboCasa, with a remarkable 11.3% improvement in OOD generalization on SimplerEnv, proving its effectiveness in breaking the vision shortcut.
@@ 2.3 Experiment 3: Catastrophic Failure in OOD Generalization
Finally, we test the generalization capability by training on the high-quality BridgeDataV2 dataset (Walke et al., 2023) (diverse, in-the-wild scenes) and evaluating on SimplerEnv (Li et al., 2024c) (simulation, OOD). During training on BridgeDataV2, the vision-only model achieves an action loss of 0.13, comparable to the full language-conditioned model’s loss of 0.08 (as shown in Figure 2(b)). This indicates that even in diverse, in-the-wild scenarios, the model can still identify visual shortcuts (e.g., specific lighting or background features mapping to specific actions) to minimize the training objective without truly grounding the language instructions.
@@ 2.4 Theoretical Insight: Information Collapse
3 Method: BayesianVLA
In this section, we introduce BayesianVLA, a framework designed to mitigate the vision shortcut in VLA models. We present the overall framework of BayesianVLA in Figure 3. We first formalize the problem through a Bayesian lens (Section 3.1), deriving an objective that maximizes the mutual information between actions and instructions. We then present our architecture, which uses Latent Action Queries to instantiate this decomposition (Section 3.2), and detail our dual-branch training strategy (Section 3.3).
Figure 3: The framework of BayesianVLA. The framework employs a dual-branch architecture with shared VLM weights. The Priori Branch (left) processes $v,\mathcal{Q},\ell]$ with causal masking to learn the vision-only prior $p(a\mid v)$. The Posteriori Branch (right) processes $v,\ell,\mathcal{Q}]$ to learn the full policy $\pi(a\mid v,\ell)$. Latent Action Queries $\mathcal{Q}$ serve as a bottleneck interface, and the LLR objective (in Eq. 7) encourages the model to maximize the information between actions and instructions. At inference, only the Posteriori Branch is used, incurring no additional computational overhead.
@@ 3.1 Objective Formulation
The detailed derivation is provided in Appendix A. This formulation requires us to simultaneously model the posterior $\pi(a\mid v,\ell)$ and the prior $p(a\mid v)$. In the following sections, we describe how BayesianVLA efficiently instantiates these two distributions using a shared architecture with Latent Action Queries.
@@ 4 Experiment
To comprehensively evaluate the effectiveness of BayesianVLA, we conducted extensive experiments on two simulation benchmarks: SimplerEnv, and RoboCasa. Our training pipeline is built upon the StarVLA framework (starVLA, 2025), distributed across 16 NVIDIA H100 GPUs, and strictly follows its default training protocols to ensure fair comparison. In our experiments, BayesianVLA is instantiated on the QwenGR00T architecture from StarVLA. Broader simulation benchmarks and real-world robot experiments are currently underway. We employ the AdamW optimizer (Loshchilov and Hutter, 2017) initialized with a learning rate of 1e-5 and a cosine annealing schedule. System-level optimizations include DeepSpeed ZeRO-2 (Rasley et al., 2020), gradient clipping at a norm of 1.0, and no gradient accumulation. All baseline performance metrics are obtained from their original papers or other peer-reviewed publications.
@@
Method Put Spoon on Towel Put Carrot on Plate Stack Green Block on Yellow Block Put Eggplant in Yellow Basket Average RT-1-X (O’Neill et al., 2024) 0.0 4.2 0.0 0.0 1.1 Octo-Small (Team et al., 2024) 41.7 8.2 0.0 56.7 26.7 OpenVLA (Kim et al., 2024) 4.2 0.0 0.0 12.5 4.2 OpenVLA-OFT (Kim et al., 2025) 12.5 4.2 4.2 72.5 23.4 RoboVLM (Li et al., 2024b) 50.0 37.5 0.0 83.3 42.7 CogACT (Li et al., 2024a) 71.7 50.8 15.0 67.5 51.3 SpatialVLA (Qu et al., 2025) 20.8 20.8 25.0 70.8 34.4 TraceVLA (Zheng et al., 2025b) 12.5 16.6 16.6 65.0 27.7 VideoVLA (Shen et al., 2025) 75.0 20.8 45.8 70.8 53.1 $\pi_{0}$ (Black et al., 2024) 29.2 62.5 29.2 91.6 53.1 $\pi_{0.5}$ (Black et al., 2025) 49.3 64.7 44.7 69.7 57.1 Isaac-GR00T-N1.6-Bridge (GEAR-Team et al., 2025) 64.5 65.5 5.5 93.0 57.1 QwenGR00T (Baseline) + Qwen3-VL-4B (starVLA, 2025) 87.5 50.0 29.2 64.2 55.2 BayesianVLA + Qwen3-VL-4B 89.6 63.8 33.3 79.2 66.5
@@ 4.1 Experiments on SimplerEnv
We utilize two large-scale subsets from the Open X-Embodiment (OXE) dataset: BridgeDataV2 (Walke et al., 2023) and Fractal (Brohan et al., 2022). The model is fine-tuned for 40k steps on a cluster of 16 GPUs (batch size 16 per device). This benchmark includes four manipulation tasks: “Put spoon on towel”, “Put carrot on plate”, “Stack green cube on yellow cube”, and “Put eggplant in yellow basket”. For each task, we evaluate the VLA policies using the official evaluation scripts provided by the SimplerEnv repository (Li et al., 2024c). To mitigate the effects of randomness, we run 480 independent trials and report the average performance (Avg@480).
@@
Task Isaac-GR00T N1.6 QwenGR00T +Qwen3VL QwenPI +Qwen3VL QwenOFT +Qwen3VL QwenFAST +Qwen3VL VisionOnly QwenGR00T BayesianVLA +Qwen3VL
@@
The results are summarized in Table 1. BayesianVLA consistently outperforms comparison baselines, achieving a state-of-the-art average success rate of 66.5%. Notably, compared to the direct baseline QwenGR00T (55.2%) built on the same StarVLA framework, our method delivers an absolute improvement of 11.3%, validating that the performance gain stems from our proposed Bayesian decomposition rather than the base architecture. Significant improvements are observed in tasks requiring precise object identification and manipulation, such as “Put Carrot on Plate” (+13.6%) and “Put Eggplant in Yellow Basket” (+15.0%). Furthermore, BayesianVLA surpasses other recent strong competitors, including the flow-matching-based $\pi_{0.5}$ (57.1%) and the dual-system Isaac-GR00T-N1.6 (57.1%). These results confirm that by explicitly optimizing the mutual information between language and action, BayesianVLA effectively mitigates the vision shortcut. Fundamentally, this validates that our approach prevents the policy from collapsing into a spurious vision-only prior $p(a|v)$ caused by dataset determinism, and instead compels the model to learn the true causal dependency of actions on language instructions.
@@ 4.2 Experiments on RoboCasa
The quantitative results on RoboCasa are presented in Table 2. Consistent with the empirical evidence presented in our motivation (Section 2), the VisionOnly baseline achieves a surprisingly high success rate of 44.7%, lagging only slightly behind the standard QwenGR00T baseline (47.8%). This observation reconfirms the prevalence of the vision shortcut in this benchmark, suggesting that a significant portion of tasks can be solved by relying solely on visual cues. However, BayesianVLA breaks this performance ceiling, achieving a state-of-the-art average success rate of 50.4% and surpassing all competing baselines, including QwenOFT (48.8%) and Isaac-GR00T (47.6%). Crucially, our method demonstrates substantial gains in tasks where the vision-only policy falters. For instance, in the task “PnP Novel From Placemat To Plate”, BayesianVLA achieves 70.0% success, significantly outperforming both the VisionOnly baseline (34.0%) and the standard QwenGR00T (48.0%). These results indicate that maximizing the LLR objective successfully forces the policy to extract and utilize task-specifying information from language, rather than settling for local optima based on visual shortcuts.
@@ 4.3 Preservation of General Capabilities
Figure 4: Qualitative comparison of general multimodal reasoning. We present a case where the model is asked to solve a mathematical problem. The standard VLA baseline (QwenGR00T) suffers from catastrophic forgetting; while the text before the comma implies “differentiating all terms together”, the subsequent output degenerates into repetitive and meaningless gibberish (bottom right). In contrast, BayesianVLA (top right) retains the VLM’s original reasoning and language generation capabilities (left), successfully solving the problem.
A prevalent view is that fine-tuning VLMs for robotic action generation (VLA training) diminishes the model’s foundational reasoning and multimodal understanding, potentially leading to a loss of general conversational abilities (Zhou et al., 2025; Xu et al., 2025; Hancock et al., 2025; Yu et al., 2026). ChatVLA (Zhou et al., 2025) attributes this to two factors: spurious forgetting, where robot training overwrites crucial visual-text alignments, and task interference, where competing control and understanding tasks degrade performance when trained jointly. We observe similar degradation in our baseline: as shown in Figure 4, the standard QwenGR00T model loses its ability to converse coherently even when prompted with pure text inputs. In contrast, BayesianVLA remarkably preserves these capabilities when queried with language instructions.
@@
It is crucial to note a nuanced distinction: while BayesianVLA preserves normal text-only conversational ability, its general vision-language conversation (image + text inputs) can still degrade after VLA training. We hypothesize this is mainly because, in our training setup, the vision tower (and the multimodal fusion/projection layers) must be adapted for control, which can shift visual representations away from the pre-trained vision–language alignment manifold. Such specialization is expected for an embodied agent optimized for action. Importantly, BayesianVLA helps prevent the text-only language behavior of the backbone from collapsing. In the baseline, the vision shortcut can render instruction tokens effectively redundant for control, weakening the training signal that encourages meaningful language processing and increasing drift in shared parameters, which manifests as failures even on pure text queries (Fig. 4 and Fig. 5). Conversely, our method’s LLR objective enforces a strong dependency on language. This acts as a regularizer that maintains the functional utility of instruction tokens, thereby preserving the backbone’s text-only conversational ability even as the visual modality is specialized for control. This preservation is of significant practical value: it ensures that the VLM backbone does not degenerate into a shallow feature mapper. By retaining its linguistic core, the agent preserves the potential for higher-level reasoning and generalization to novel instructions, which are the primary motivations for employing Foundation Models in robotics.

### v4

# LangForce: Bayesian Decomposition of Vision Language Action Models via Latent Action Queries

@@ Abstract
Vision-Language-Action (VLA) models have shown promise in robot manipulation but often struggle to generalize to new instructions or complex multi-task scenarios. We identify a critical pathology in current training paradigms where goal-driven data collection creates a dataset bias. In such datasets, language instructions are highly predictable from visual observations alone, causing the conditional mutual information between instructions and actions to vanish, a phenomenon we term Information Collapse. Consequently, models degenerate into vision-only policies that ignore language constraints and fail in out-of-distribution (OOD) settings. To address this, we propose LangForce, a novel framework that enforces instruction following via Bayesian decomposition. By introducing learnable Latent Action Queries, we construct a dual-branch architecture to estimate both a vision-only prior $p(a\mid v)$ and a language-conditioned posterior $\pi(a\mid v,\ell)$. We then optimize the policy to maximize the conditional Pointwise Mutual Information (PMI) between actions and instructions. This objective effectively penalizes the vision shortcut and rewards actions that explicitly explain the language command. Without requiring new data, LangForce significantly improves generalization. Extensive experiments across on SimplerEnv and RoboCasa demonstrate substantial gains, including an 8.8% improvement on the challenging OOD SimplerEnv benchmark, validating the ability of our approach to robustly ground language in action.
LangForce: Bayesian Decomposition of Vision Language Action Models via Latent Action Queries
Shijie Lian1,2,thanks: Equal contribution Bin Yu2,4,11footnotemark: 1 Xiaopeng Lin2,5,11footnotemark: 1 Laurence T. Yang6,1,thanks: Corresponding author Zhaolong Shen2,7 Changti Wu2,8 Yuzhuo Miao2,4 Cong Huang2,3 Kai Chen2,3,9,22footnotemark: 2 1HUST 2ZGCA 3ZGCI 4HIT 5HKUST(GZ) 6ZZU 7BUAA 8ECNU 9DeepCybo
@@ 1 Introduction
While demonstrating strong performance in in-distribution settings, current VLA models still face challenges in generalizing to novel instructions or complex multi-task scenarios, particularly in out-of-distribution (OOD) environments (Xing et al., 2025; Xu et al., 2023, 2025). This limitation is especially pronounced during post-training, where fine-tuning on narrow, task-specific datasets can lead to catastrophic forgetting of the VLM’s general capabilities and impair its ability to generalize to new tasks (Xing et al., 2025; Zhou et al., 2025). We hypothesize that this fragility is exacerbated by a prevalent bias in current robotic datasets. Most robotic datasets are collected in a goal-driven manner, where a human operator performs a specific task repeatedly in a fixed scene. In such datasets, the mapping from visual scene $v$ to language instruction $\ell$ is nearly injective: seeing a cabinet in the scene almost invariably implies the task “open the cabinet,” while seeing a bottle implies “pick up the bottle.” This deterministic coupling results in a sharp conditional distribution $p(\ell\mid v)$.
@@
In other words, the model effectively ignores the language instruction, learning a “vision shortcut” that fails whenever the task is ambiguous or the environment changes. We provide a rigorous probabilistic analysis of why this collapse occurs in Appendix A.
To address this, we propose LangForce, a novel framework that explicitly enforces instruction following via Bayesian decomposition. Our key insight is to maximize the conditional Pointwise Mutual Information (PMI) between actions and instructions, which is equivalent to maximizing the log-likelihood ratio (LLR): $\log p(\ell\mid a,v)-\log p(\ell\mid v)$. This objective penalizes the vision shortcut by requiring the action $a$ to provide additional information about $\ell$ that cannot be inferred from $v$ alone.
We instantiate this framework by introducing Latent Action Queries—a set of learnable tokens injected into the VLM. These queries serve a dual purpose: they act as a bottleneck to extract action-relevant features for a downstream Diffusion Transformer (DiT) policy, and they enable a dual-branch training strategy. In the Priori Branch, queries attend only to vision to learn $p(a\mid v)$; in the Posteriori Branch, they attend to both vision and language to learn $\pi(a\mid v,\ell)$. By optimizing the LLR between these branches, LangForce learns to ground language robustly without requiring new data.
@@
    We propose LangForce, a method that leverages Latent Action Queries and a dual-branch Bayesian objective to recover the true language-conditioned policy from biased data.
@@
    We demonstrate that LangForce achieves state-of-the-art performance on SimplerEnv and RoboCasa, with a remarkable 8.8% improvement in OOD generalization on SimplerEnv, proving its effectiveness in breaking the vision shortcut.
@@ 2.3 Experiment 3: Catastrophic Failure in OOD Generalization
Finally, we test the generalization capability by training on the high-quality BridgeDataV2 dataset (Walke et al., 2023) (diverse, in-the-wild scenes) and evaluating on SimplerEnv (Li et al., 2024d) (simulation, OOD). During training on BridgeDataV2, the vision-only model achieves an action loss of 0.13, comparable to the full language-conditioned model’s loss of 0.08 (as shown in Figure 2(b)). This indicates that even in diverse, in-the-wild scenarios, the model can still identify visual shortcuts (e.g., specific lighting or background features mapping to specific actions) to minimize the training objective without truly grounding the language instructions.
@@ 2.4 Theoretical Insight: Information Collapse
3 Method: LangForce
In this section, we introduce LangForce, a framework designed to mitigate the vision shortcut in VLA models. We present the overall framework of LangForce in Figure 3. We first formalize the problem through a Bayesian lens (Section 3.1), deriving an objective that maximizes the mutual information between actions and instructions. We then present our architecture, which uses Latent Action Queries to instantiate this decomposition (Section 3.2), and detail our dual-branch training strategy (Section 3.3).
Figure 3: The framework of LangForce. The framework employs a dual-branch architecture with shared VLM weights. The Priori Branch (left) processes $v,\mathcal{Q},\ell]$ with causal masking to learn the vision-only prior $p(a\mid v)$. The Posteriori Branch (right) processes $v,\ell,\mathcal{Q}]$ to learn the full policy $\pi(a\mid v,\ell)$. Latent Action Queries $\mathcal{Q}$ serve as a bottleneck interface, and the LLR objective (in Eq. 7) encourages the model to maximize the information between actions and instructions. At inference, only the Posteriori Branch is used, incurring no additional computational overhead.
@@ 3.1 Objective Formulation
The detailed derivation is provided in Appendix B. This formulation requires us to simultaneously model the posterior $\pi(a\mid v,\ell)$ and the prior $p(a\mid v)$. In the following sections, we describe how LangForce efficiently instantiates these two distributions using a shared architecture with Latent Action Queries.
@@ 4 Experiment
To comprehensively evaluate the effectiveness of LangForce, we conducted extensive experiments on two simulation benchmarks: SimplerEnv, and RoboCasa. Our training pipeline is built upon the StarVLA framework (starVLA, 2025), distributed across 16 NVIDIA H100 GPUs, and strictly follows its default training protocols to ensure fair comparison. In our experiments, LangForce is instantiated on the QwenGR00T architecture from StarVLA. Broader simulation benchmarks and real-world robot experiments are currently underway. We employ the AdamW optimizer (Loshchilov and Hutter, 2017) initialized with a learning rate of 1e-5 and a cosine annealing schedule. System-level optimizations include DeepSpeed ZeRO-2 (Rasley et al., 2020), gradient clipping at a norm of 1.0, and no gradient accumulation. All baseline performance metrics are obtained from their original papers or other peer-reviewed publications.
@@
Method Put Spoon on Towel Put Carrot on Plate Stack Green Block on Yellow Block Put Eggplant in Yellow Basket Average RT-1-X (O’Neill et al., 2024) 0.0 4.2 0.0 0.0 1.1 Octo-Small (Team et al., 2024) 41.7 8.2 0.0 56.7 26.7 OpenVLA (Kim et al., 2024) 4.2 0.0 0.0 12.5 4.2 OpenVLA-OFT (Kim et al., 2025) 12.5 4.2 4.2 72.5 23.4 RoboVLM (Li et al., 2024c) 50.0 37.5 0.0 83.3 42.7 CogACT (Li et al., 2024b) 71.7 50.8 15.0 67.5 51.3 SpatialVLA (Qu et al., 2025) 20.8 20.8 25.0 70.8 34.4 TraceVLA (Zheng et al., 2025b) 12.5 16.6 16.6 65.0 27.7 VideoVLA (Shen et al., 2025) 75.0 20.8 45.8 70.8 53.1 $\pi_{0}$ (Black et al., 2024) 29.2 62.5 29.2 91.6 53.1 $\pi_{0.5}$ (Black et al., 2025) 49.3 64.7 44.7 69.7 57.1 Isaac-GR00T-N1.6-Bridge (GEAR-Team et al., 2025) 64.5 65.5 5.5 93.0 57.1 QwenGR00T (Baseline) + Qwen3-VL-4B (starVLA, 2025) 87.5 50.0 29.2 64.2 57.7 LangForce + Qwen3-VL-4B 89.6 63.8 33.3 79.2 66.5
@@ 4.1 Experiments on SimplerEnv
We utilize two large-scale subsets from the Open X-Embodiment (OXE) dataset: BridgeDataV2 (Walke et al., 2023) and Fractal (Brohan et al., 2022). The model is fine-tuned for 40k steps on a cluster of 16 GPUs (batch size 16 per device). This benchmark includes four manipulation tasks: “Put spoon on towel”, “Put carrot on plate”, “Stack green cube on yellow cube”, and “Put eggplant in yellow basket”. For each task, we evaluate the VLA policies using the official evaluation scripts provided by the SimplerEnv repository (Li et al., 2024d). To mitigate the effects of randomness, we run 480 independent trials and report the average performance (Avg@480).
@@
Task Isaac-GR00T N1.6 QwenGR00T +Qwen3VL QwenPI +Qwen3VL QwenOFT +Qwen3VL QwenFAST +Qwen3VL VisionOnly QwenGR00T LangForce +Qwen3VL
@@
The results are summarized in Table 1. LangForce consistently outperforms comparison baselines, achieving a state-of-the-art average success rate of 66.5%. Notably, compared to the direct baseline QwenGR00T (55.2%) built on the same StarVLA framework, our method delivers an absolute improvement of 8.8%, validating that the performance gain stems from our proposed Bayesian decomposition rather than the base architecture. Significant improvements are observed in tasks requiring precise object identification and manipulation, such as “Put Carrot on Plate” (+13.6%) and “Put Eggplant in Yellow Basket” (+15.0%). Furthermore, LangForce surpasses other recent strong competitors, including the flow-matching-based $\pi_{0.5}$ (57.1%) and the dual-system Isaac-GR00T-N1.6 (57.1%). These results confirm that by explicitly optimizing the mutual information between language and action, LangForce effectively mitigates the vision shortcut. Fundamentally, this validates that our approach prevents the policy from collapsing into a spurious vision-only prior $p(a|v)$ caused by dataset determinism, and instead compels the model to learn the true causal dependency of actions on language instructions.
@@ 4.2 Experiments on RoboCasa
The quantitative results on RoboCasa are presented in Table 2. Consistent with the empirical evidence presented in our motivation (Section 2), the VisionOnly baseline achieves a surprisingly high success rate of 44.7%, lagging only slightly behind the standard QwenGR00T baseline (47.8%). This observation reconfirms the prevalence of the vision shortcut in this benchmark, suggesting that a significant portion of tasks can be solved by relying solely on visual cues. However, LangForce breaks this performance ceiling, achieving a state-of-the-art average success rate of 50.4% and surpassing all competing baselines, including QwenOFT (48.8%) and Isaac-GR00T (47.6%). Crucially, our method demonstrates substantial gains in tasks where the vision-only policy falters. For instance, in the task “PnP Novel From Placemat To Plate”, LangForce achieves 70.0% success, significantly outperforming both the VisionOnly baseline (34.0%) and the standard QwenGR00T (48.0%). These results indicate that maximizing the LLR objective successfully forces the policy to extract and utilize task-specifying information from language, rather than settling for local optima based on visual shortcuts.
@@ 4.3 Preservation of General Capabilities
Figure 4: Qualitative comparison of general multimodal reasoning. We present a case where the model is asked to solve a mathematical problem. The standard VLA baseline (QwenGR00T) suffers from catastrophic forgetting; while the text before the comma implies “differentiating all terms together”, the subsequent output degenerates into repetitive and meaningless gibberish (bottom right). In contrast, LangForce (top right) retains the VLM’s original reasoning and language generation capabilities (left), successfully solving the problem.
A prevalent view is that fine-tuning VLMs for robotic action generation (VLA training) diminishes the model’s foundational reasoning and multimodal understanding, potentially leading to a loss of general conversational abilities (Zhou et al., 2025; Xu et al., 2025; Hancock et al., 2025; Yu et al., 2026). ChatVLA (Zhou et al., 2025) attributes this to two factors: spurious forgetting, where robot training overwrites crucial visual-text alignments, and task interference, where competing control and understanding tasks degrade performance when trained jointly. We observe similar degradation in our baseline: as shown in Figure 4, the standard QwenGR00T model loses its ability to converse coherently even when prompted with pure text inputs. In contrast, LangForce remarkably preserves these capabilities when queried with language instructions.
@@
It is crucial to note a nuanced distinction: while LangForce preserves normal text-only conversational ability, its general vision-language conversation (image + text inputs) can still degrade after VLA training. We hypothesize this is mainly because, in our training setup, the vision tower (and the multimodal fusion/projection layers) must be adapted for control, which can shift visual representations away from the pre-trained vision–language alignment manifold. Such specialization is expected for an embodied agent optimized for action. Importantly, LangForce helps prevent the text-only language behavior of the backbone from collapsing. In the baseline, the vision shortcut can render instruction tokens effectively redundant for control, weakening the training signal that encourages meaningful language processing and increasing drift in shared parameters, which manifests as failures even on pure text queries (Fig. 4 and Fig. 5). Conversely, our method’s LLR objective enforces a strong dependency on language. This acts as a regularizer that maintains the functional utility of instruction tokens, thereby preserving the backbone’s text-only conversational ability even as the visual modality is specialized for control. This preservation is of significant practical value: it ensures that the VLM backbone does not degenerate into a shallow feature mapper. By retaining its linguistic core, the agent preserves the potential for higher-level reasoning and generalization to novel instructions, which are the primary motivations for employing Foundation Models in robotics.